### v1

# Edge-weighted Online Stochastic Matching Under Jaillet-Lu LP

Edge-weighted Online Stochastic Matching Under Jaillet-Lu LP
Shuyi Yan BARC, University of Copenhagen. Supported by VILLUM Foundation Grant 54451. Email: shya@di.ku.dk.
The online stochastic matching problem was introduced by Feldman et al. [7], together with the $(1-\frac{1}{e})$-competitive Suggested Matching algorithm. In the most general edge-weighted setting, this ratio has not been improved for more than one decade, until recently Yan [21] beat the $1-\frac{1}{e}$ bound and Qiu et al. [19] further improved the ratio to $0.650$. Both of these works measure the online competitiveness against the offline LP relaxation introduced by Jaillet and Lu [14]. This LP has also played an important role in other settings since it is a natural choice for two-choices online algorithms.
In this paper, we propose an upper bound of $0.663$ and a lower bound of $0.662$ for edge-weighted online stochastic matching under Jaillet-Lu LP. First, we propose a hard instance and prove that the optimal online algorithm for this instance only has a competitive ratio $<0.663$. Then, we show that a near-optimal algorithm for this instance can be generalized to work on all instances and achieve a competitive ratio $>0.662$. It indicates that more powerful LPs are necessary if we want to further improve the ratio by $0.001$.
1 Introduction
The online bipartite matching problem was introduced by Karp, Vazirani, and Vazirani [16], motivated by applications such as online advertising. In this problem, we are given a bipartite graph. Advertisers are modeled as offline vertices on one side of the graph, which are known at the beginning. Impressions (user queries) are modeled as online vertices on the other side, which arrive one by one. The weights on the edges indicate how much the advertisers are interested in the impressions. When an online vertex arrives, the online algorithm sees its incident edges, and must immediately and irrevocably decide how to match it. Karp et al. [16] proposed the celebrated Ranking algorithm, which achieves a competitive ratio of $1-\frac{1}{e}\approx 0.632$ in the unweighted setting. This is optimal in the worst case model, where we assume the online algorithm has no prior information about the graph.
However, in real-world scenarios, the algorithm often knows some prior information. For example, online advertising platforms can use historic data to predict user behaviors in the future. Feldman, Mehta, Mirrokni, and Muthukrishnan [7] introduced the online stochastic matching model, where we assume all online vertices are drawn independently from an identical distribution which is known by the algorithm in the beginning. Feldman et al. [7] proposed the Suggested Matching algorithm which is $(1-\frac{1}{e})$-competitive even in the most difficult edge-weighted setting, and then beat the $1-\frac{1}{e}$ bound in the unweighted setting under the integral assumption [^1] It means that the expected number of online vertices of each type is an integer.. Further works removed the integral assumption [18, 14, 11, 13, 20] and beat the $1-\frac{1}{e}$ bound in the vertex-weighted setting [11, 13, 20] and the free-disposal [^2] It allows us to match an offline vertex many times and only count the most valuable one in the end. edge-weighted setting [13].
A common framework of the online algorithms is to solve an LP to get a fractional matching which upperbounds the offline optimal solution, and then use the fractional solution to guide the online strategy. The online competitiveness is then measured against the LP relaxation. In particular, the LP proposed by Jaillet and Lu [14] is quite useful for two-choices algorithms (which most previous algorithms are) in the sense that it limits the case that an online vertex type is mostly matched to one offline vertex in the fractional solution, forcing most online vertices to have at least two choices. Under Jaillet-Lu LP, Huang, Shu, and Yan [13] provided both a lower bound and an upper bound of the same ratio 0.7060.7060.7060.706333The exact value is $1-\frac{1}{1-\ln 2}\big{(}\frac{1}{2e}-\frac{\ln 2}{e^{2}}\big{)}$. for unweighted, vertex-weighted and free-disposal edge-weighted settings.
The edge-weighted setting was known to be strictly harder than the above settings [13]. Recently, Yan [21] beat the $1-\frac{1}{e}$ bound with a competitive ratio of $0.645$ by extending the Suggested Matching algorithm to have different strategies in $3$ different time periods. Qiu, Feng, Zhou, and Wu [19] improved the ratio to $0.650$ by increasing the number of periods to $5$. Both algorithms still use the trivial strategy from Suggested Matching in more than half of the time. In contrast, instead of using several fixed strategies in different periods, the state-of-the-art Poisson Online Correlated Selection algorithm [13] for unweighted/vertex-weighted settings has a strategy which smoothly changes over time. A natural question arises: Can smoothly changed strategies help to further improve the competitive ratio in the edge-weighted setting?
On the other hand, both of [21, 19] use the Jaillet-Lu LP relaxation. We know that the provably best ratio for online algorithms under Jaillet-Lu LP is $0.706$ in the other $3$ settings. What about the edge-weighted setting?
1.1 Our Results
In this paper, we prove that for edge-weighted online stochastic matching, the optimal competitive ratio against Jaillet-Lu LP is between $0.662$ and $0.663$.
We first propose a hard instance and show that, for this instance, the optimal online algorithm can only achieve a competitive ratio $<0.663$. Then, we extract some key properties of a near-optimal matching process of this specific instance which are sufficient to guarantee a competitive ratio $>0.662$. Finally, we propose an algorithm which maintains these properties for all instances, which is therefore $0.662$-competitive for all instances. The algorithm improves the best known ratio of $0.650$ [19], and the hardness result suggests that more powerful LPs are necessary for further improvements.
Compared with all existing algorithms for edge-weighted online stochastic matching, our algorithm has two major differences: (1) Instead of only using one or several fixed strategies, we deploy a strategy that varies from time to time; (2) When an online vertex arrives, instead of only locally looking at its neighbors, our algorithm utilizes some global information about the probability distribution of remaining offline vertices.
Below, we briefly summarize our high-level ideas. For readers not familiar with this problem, detailed definitions and explanations can be found in Section 2.
Hardness.
Previous work [21] showed that, under Jaillet-Lu LP, the problem can be reduced to the case that there are only two classes of online vertex types:
  • A first-class type has only one offline neighbor;
  • A second-class type has only two offline neighbors, and it is equally matched to them.
In the hard instance, we want to have as many (arrival rates of) first-class types as possible, since they don’t have a second choice when the only neighbor has been matched. For second-class types, we want to maximize the probability that, given that one neighbor has been matched, the other neighbor has also been matched. So, if two second-class types share one common neighbor, we tend to let them share another. The above ideas give us the following instance:
  • For each offline vertex, there is a first-class online vertex type matched to it, with an arrival rate as high as possible;
  • Offline vertices are divided into pairs. For each pair, there is a second-class online vertex type matched to them.
By carefully setting the edge weights, we show that the optimal algorithm for this instance only achieves a competitive ratio $<0.663$.
Algorithm.
We modify the optimal algorithm for the hard instance to get a near-optimal algorithm $A$, which on this instance will match each edge $e$ with probability $\Gamma x_{e}$, where $\Gamma>0.662$ and $x_{e}$ is the matching fraction of $e$ in the optimal fractional matching. We want it to be generalized to work for all instances.
First-class types are easy to deal with since they have only one possible choice when arriving, so we focus on second-class types. For each second-class type, we want to match it to each neighbor at the same rate as $A$. To make this feasible, we need an upper bound of the probability that two offline vertices have both been matched at some time $t$. We show that this probability takes its maximum value in our hard instance when we “simulate” $A$, and this finally allows us to achieve a competitive ratio of $\Gamma>0.662$.
1.2 Other Related Work
In the worst case model, the Ranking algorithm [16] and its generalization [1] achieved the optimal competitive ratio of $1-\frac{1}{e}$ for unweighted and vertex-weighted matching. For free-disposal edge-weighted matching, a sequence of works [6, 8, 3] improved the competitive ratio to $0.5239$.
In the random order model [^4] The adversary decides the graph, then online vertices arrive in a uniformly random order., the currently best ratios are $0.696$ [9, 17] for unweighted matching and $0.662$ [12, 15] for vertex-weighted matching.
In the online stochastic matching model, for unweighted and vertex-weighted matching, a sequence of works [7, 2, 18, 14, 11, 13, 20] improved the competitive ratio to $0.716$ and the upper bound to $0.823$. For free-disposal edge-weighted matching, Huang et al. [13] proposed a $0.706$-competitive algorithm. Under the integral assumption, the state-of-the-art algorithms achieved competitive ratios of $0.7299$ for unweighted matching [18, 14, 4], $0.725$ for vertex-weighted matching [14], and $0.705$ for edge-weighted matching [10, 4]. An upper bound of $0.703$ [13] for (general) edge-weighted matching inferred that it is strictly harder than the above settings. For non i.i.d. arrivals, recent works [20, 5] achieved a competitive ratio of $0.69$ for unweighted and vertex-weighted matching.
2 Preliminaries
Edge-weighted Online Stochastic Matching.
An instance of edge-weighted online stochastic matching consists of:
  • A bipartite graph $G=(I,J,E)$, where $I$ is the set of online vertex types and $J$ is the set of offline vertices;
  • An arrival rate $\lambda_{i}$ for each online vertex type $i\in I$;
  • A nonnegative weight $w_{ij}$ for each edge $(i,j)\in E$.
$\Lambda=\sum_{i\in I}\lambda_{i}$ is the total number of online vertices. $\Lambda$ online vertices will arrive one by one. When an online vertex arrives, it independently draws a random type $i\in I$ with probability $\lambda_{i}/\Lambda$. Then, the online algorithm can either match it to an unmatched offline neighbor, or simply discard it. The decision has to be made immediately and irrevocably. The objective is to maximize the expected total weight of all matched edges in the end. The competitive ratio of the online algorithm is the infimum of the ratio of its expected objective to the expected objective of the offline optimal matching.
Poisson Arrival Model.
In the Poisson arrival model, instead of assuming a deterministic number of online vertices, we assume that online vertices arrive following a Poisson process with arrival rate $\Lambda$ in the time horizon $[0,1]$. An equivalent view is that each online vertex type $i\in I$ independently follows a Poisson process with arrival rate $\lambda_{i}$. This equivalency and the continuous arrival process greatly help the analyses of online stochastic matching algorithms, after Huang and Shu [11] proved that the Poisson arrival model is asymptotically equivalent to the original model. In this paper, we present our upper and lower bounds in the Poisson arrival model. We remark that the same bounds hold in the original model with an additive term of $o(1)$ with respect to $\Lambda$. We also remark that, prior works in the original model also conventionally assume a large enough $\Lambda$, or explicitly have the $o(1)$ additive term.
Jaillet-Lu LP.
The (Poisson arrival version of) Jaillet-Lu LP [14] is as follows:
maximize $\sum_{(i,j)\in E}w_{ij}x_{ij}$
subject to $\sum_{j\in J}x_{ij}\leq\lambda_{i}$ $\forall i\in I$
$\sum_{i\in I}x_{ij}\leq 1$ $\forall j\in J$
$\sum_{i\in I}\max\{2x_{ij}-\lambda_{i},0\}\leq 1-\ln 2$ $\forall j\in J$ (1)
$x_{ij}\geq 0$ $\forall(i,j)\in E$
A solution of the Jaillet-Lu LP is a fractional matching between online vertex types and offline vertices. Huang and Shu [11] proved that the optimal value of the LP, that is, the objective of the optimal fractional matching, upperbounds the expected objective of the optimal offline integral matching. In this paper, we consider the competitive ratio against the optimal value of the Jaillet-Lu LP instead of the optimal integral matching.
Without loss of generality, we assume $x_{ij}=0$ for any pair $(i,j)\notin E$. For each online vertex type $i$, define $x_{i}=\sum_{j\in J}x_{ij}$. For each offline vertex $j$, define $x_{j}=\sum_{i\in I}x_{ij}$.
Instance Reduction.
Yan [21] showed that, any instance $G$ can be efficiently transformed into another instance $G^{\prime}$ such that the optimal fractional matching (under Jaillet-Lu LP) of $G^{\prime}$ satisfies the following conditions:
  • Each offline vertex is fully matched. In other words, $x_{j}=1$ for each offline vertex $j$.
  • There are only two classes of online vertex types:
    • A first-class online vertex type $i$ only has one offline neighbor $j$, where $x_{ij}=x_{i}=\lambda_{i}$.
    • A second-class online vertex type $i$ has exactly two offline neighbors $j_{1}$ and $j_{2}$, where $x_{ij_{1}}=x_{ij_{2}}=\lambda_{i}/2$ and $x_{i}=\lambda_{i}$.
In addition, for any online algorithm which works for $G^{\prime}$, using the transformation as a preprocessing step, it works for $G$ with the same competitive ratio. So, without loss of generality, in this paper we restrict the instance to satisfy the above conditions.
If an edge is incident to a first-class (resp. second-class) online vertex type, we call it a first-class (resp. second-class) edge. For each offline vertex $j$, let $y_{j}$ denote the fraction that $j$ is matched to first-class vertices, i.e.
$y_{j}=\sum_{i:\text{first-class}}x_{ij}.$
Then Equation 1 is saying that $y_{j}\leq 1-\ln 2$.
3 Hardness
We define our hard instance $G$ with $\Lambda=2$ as follows. There are two offline vertices and a second-class online vertex type, with arrival rate $2\ln 2$, connected to both offline vertices with unit edge weights. In addition, each offline vertex is connected to a first-class online vertex type with arrival rate $1-\ln 2$, with some edge weight $k\geq 1$. We will choose $k$ later. See Figure 1 for an illustration. We note that we can also get hard instances with arbitrarily large $\Lambda$ by simply copying $G$ many times.
Figure 1: The hard instance $G$. The number next to each online vertex type is the arrival rate. The number on each edge is the edge weight.
It’s easy to check that in the optimal Jaillet-Lu LP solution, all vertices in $G$ will be fully matched, which gives us the following fact.
Fact 1.
The optimal value of Jaillet-Lu LP on $G$ is $2\ln 2+(2-\ln 2)k$.
Now we are going to find the optimal online algorithm on $G$. Clearly, since $k\geq 1$, when a first-class online vertex arrives, we shall always match it if we can. What about a second-class online vertex? Intuitively, the decision should depend on $k$ and the current time $t$. When $t$ is small but $k$ is large, we may want to discard this online vertex even if it has unmatched neighbors, to preserve the chance of matching a second-class vertex in the future. Also notice that, the number of unmatched neighbors may also affect our decision, since it affects the chance that the offline vertex can be matched with another second-class online vertex in the future.
Let’s first look at an easier case where we don’t need to consider the number of unmatched neighbors. Let $G^{\prime}$ be the remaining part of $G$ after one offline vertex is matched. See Figure 2.
Figure 2: The sub-instance $G^{\prime}$. The online vertex type with no unmatched neighbor can be ignored.
Algorithm 1 Optimal Algorithm on $G^{\prime}$
1:When a first-class online vertex arrives, match it to the neighbor if the neighbor is unmatched.
2:When a second-class online vertex arrives at time $t$, match it to the neighbor if the neighbor is unmatched and $t>t_{1}$.
Lemma 2.
For any $k\geq 1$, there exists $0\leq t_{1}\leq 1$ such that Algorithm 1 is the optimal algorithm on $G^{\prime}$.
Proof.
Let $\textsc{Opt}(G^{\prime},t)$ denote the expected objective of the optimal online algorithm running on $G^{\prime}$ in the time interval $[t,1]$. Clearly, $\textsc{Opt}(G^{\prime},t)$ is at most $k$ and decreasing on $t$. Consider what the optimal algorithm should do when an online vertex arrives at time $t$ and the offline vertex is still unmatched. If the algorithm decides to discard the vertex, it will gain $\textsc{Opt}(G^{\prime},t)$ in the future by definition. Since $\textsc{Opt}(G^{\prime},t)\leq k$, the algorithm can always decide to match the vertex if it is first-class. If it is second-class, the algorithm can match it if $\textsc{Opt}(G^{\prime},t)\leq 1$. Since $\textsc{Opt}(G^{\prime},t)$ is decreasing, $t_{1}$ can be set as the boundary where $\textsc{Opt}(G^{\prime},t)$ starts going below $1$. ∎
Clearly, the optimal algorithm on $G$ will degenerate to Algorithm 1 after one offline vertex is matched. When both offline vertices are unmatched, it needs another parameter $t_{0}$.
Algorithm 2 Optimal Algorithm on $G$
1:When a first-class online vertex arrives, match it to the neighbor if the neighbor is unmatched.
2:When a second-class online vertex with exactly one unmatched neighbor arrives at time $t$, match it to the neighbor if $t>t_{1}$.
3:When a second-class online vertex with two unmatched neighbors arrives at time $t$, match it to an arbitrary neighbor if $t>t_{0}$.
Lemma 3.
For any $k\geq 1$, there exists $0\leq t_{0}\leq t_{1}\leq 1$ such that Algorithm 2 is the optimal algorithm on $G$.
Proof.
When one of the offline vertices has been matched, $G$ will degenerate to $G^{\prime}$, and Algorithm 2 will also degenerate to Algorithm 1. Lemma 2 shows that Algorithm 1 is optimal in this case for some $t_{1}$.
Then we analyze the strategy of the optimal online algorithm when both offline vertices are unmatched. Let $\textsc{Opt}(G,t)$ denote the expected objective of the optimal online algorithm running on $G$ in the time interval $[t,1]$. Notice that $\textsc{Opt}(G,t)\leq 2\textsc{Opt}(G^{\prime},t)$, since $2\textsc{Opt}(G^{\prime},t)$ can be viewed as the optimal expected objective in the case that a second-class online vertex is allowed to be simultaneously matched to both neighbors on $G$.
When a first-class online vertex arrives, if the algorithm decides to match it, the expected objective in the end will be $\textsc{Opt}(G^{\prime},t)+k\geq 2\textsc{Opt}(G^{\prime},t)\geq\textsc{Opt}(G,t)$, so the algorithm can always match it.
When a second-class online vertex arrives, the algorithm can match it if $\textsc{Opt}(G^{\prime},t)+1\geq\textsc{Opt}(G,k)$, and it doesn’t matter which neighbor is matched by symmetry. When $t>t_{1}$, we know $\textsc{Opt}(G^{\prime},t)\leq 1$, then $\textsc{Opt}(G^{\prime},t)+1\geq 2\textsc{Opt}(G^{\prime},t)\geq\textsc{Opt}(G,t)$. So the algorithm can always match it when $t>t_{1}$. Then, it remains to show that $\textsc{Opt}(G,t)-\textsc{Opt}(G^{\prime},t)$ is decreasing in $[0,t_{1}]$. If this is true, then we can set $t_{0}$ as the boundary where $\textsc{Opt}(G,t)-\textsc{Opt}(G^{\prime},t)$ starts going below $1$.
Denote the two offline vertices by $u$ and $v$. The basic idea is that, by the symmetry between $u$ and $v$, we can introduce some constraints to the optimal algorithm to force $u$ to contribute exactly $\textsc{Opt}(G^{\prime},t)$ to the expected objective. Then, $v$ must contribute $\textsc{Opt}(G,t)-\textsc{Opt}(G^{\prime},t)$. We can show that for $t^{\prime}<t$, there is an algorithm for $t^{\prime}$ which simulates the optimal algorithm for $t$ on $v$ while still having optimal performance on $u$.
Note that, by Lemma 2, when $t<t_{1}$, the optimal algorithm on $G^{\prime}$ will not match any second-class vertex. So, we can force the optimal algorithm on $G$ to never match $u$ before time $t_{1}$. On the other hand, since the algorithm will decide to match a second-class online vertex after $t_{1}$, we can also force it to match to $u$ as long as possible. With these two constraints, the algorithm still achieve an expected objective of $\textsc{Opt}(G,t)$, and on $u$ it actually acts the same as the optimal algorithm on $G^{\prime}$, which means $u$ contributes exactly $\textsc{Opt}(G^{\prime},t)$ to the expected objective.
Now consider an algorithm for some $t^{\prime}<t$ which satisfies the two constraints and always match a first-class vertex if possible. It’s straight-forward to see that $u$ must contribute $\textsc{Opt}(G^{\prime},t^{\prime})$. On the other hand, at any time in $[t,1]$, the probability that $u$ has been matched in our algorithm is at least the probability that $u$ has been matched in the optimal algorithm for $t$. So, we can simply do not match $v$ in $[t^{\prime},t)$, and then in $[t,1]$ match $v$ in the same way as the optimal algorithm for $t$. Then the expected objective of our algorithm is $\textsc{Opt}(G^{\prime},t^{\prime})+\textsc{Opt}(G,t)-\textsc{Opt}(G^{\prime},t)$, which means $\textsc{Opt}(G,t^{\prime})-\textsc{Opt}(G^{\prime},t^{\prime})\geq\textsc{Opt}(G,t)-\textsc{Opt}(G^{\prime},t)$. In other words, $\textsc{Opt}(G,t)-\textsc{Opt}(G^{\prime},t)$ is decreasing in $[0,t_{1}]$ as desired. ∎
Combining 1 and 3, we can get our hardness result.
Theorem 4.
For $k\approx 3.40216$, the optimal online algorithm has a competitive ratio $\approx 0.66275<0.663$ against the optimal value of the Jaillet-Lu LP on $G$.
Proof.
For any $k,t_{0},t_{1}$, let $\textsc{Alg}(k,t_{0},t_{1})$ denote the expected objective of Algorithm 2. According to our matching strategy, $\textsc{Alg}(k,t_{0},t_{1})$ can be calculated using straightforward calculus. However, the closed-form expression is super long, so we omit it here. See Section A.1 for computational details.
By 1, the optimal LP value is $2\ln 2+(2-\ln 2)k$. So the competitive ratio is $\frac{\textsc{Alg}(k,t_{0},t_{1})}{2\ln 2+(2-\ln 2)k}$. For $k\approx 3.40216$, the ratio achieves its maximum value $\approx 0.66275$ at $t_{0}\approx 0.12437$ and $t_{1}\approx 0.29539$. ∎
4 Algorithm
In this section, we are going to generalize a special algorithm on our hard instance $G$ to get the same competitive ratio on all possible instances. The basic idea is, for each first-class (resp. second-class) edge, we want to match it at the same rate as we match a first-class (resp. second-class) edge of $G$. Note that this also fixes the match rate of each offline vertex.
Is it possible? Consider any pair of offline vertices. Without loss of generality, we can assume there is a second-class online vertex type (with an arrival rate of almost 00) adjacent to them. Intuitively, to make sure that we can match them at the desired rate, we wish that these two offline vertices are matched in a very negatively correlated way, so that there is enough probability that the second-class online vertex has at least one unmatched neighbor. In other words, a necessary condition is, $G$ must be the instance that the pair of offline vertices are matched in the most positively correlated way (with respect to the match rate of each first-class/second-class edge, which is decided by the algorithm).
It’s not hard to see that in Algorithm 2, unfortunately, the two offline vertices are matched in the most negatively correlated way during time $[t_{0},t_{1}]$: When one of them is matched, the other must not be matched by a second-class online vertex. However, during other times, they are matched in the most positively correlated way, as we will see later. [^5] During time $[0,t_{0}]$, they are matched independently, which is also the most positively correlated way given the match rate of each first-class/second-class edge, since the match rate of each second-class edge is 00. So, if we restrict $t_{0}=t_{1}$ to remove the middle phase, we will have the desired property.
Algorithm 3 Restricted version of Algorithm 2
1:When a first-class online vertex arrives, match it to the neighbor if the neighbor is unmatched.
2:When a second-class online vertex with at least one unmatched neighbor arrives at time $t>t_{0}$, match it to a uniformly random unmatched neighbor.
If we pick the optimal $t_{0}$, Algorithm 3 is slightly worse than Algorithm 2.
Lemma 5.
For $t_{0}\approx 0.14753$, Algorithm 3 matches each first-class edge (in $G$) with probability $\Gamma(1-\ln 2)$ and each second-class edge with probability $\Gamma\ln 2$, where $\Gamma\approx 0.66217>0.662$.
Proof.
Given that we have analyzed the entire matching process of Algorithm 2 in Section A.1, it is straightforward to verify the lemma. See Section A.2 for computational details. ∎
There is another issue for generalizing Algorithm 3. In the hard instance $G$, every offline vertex $j$ satisfies $y_{j}=1-\ln 2$, i.e. it is maximally matched to first-class vertices. However, in general we may have various $y_{j}\in[0,1-\ln 2]$. As we have different strategies for first-class and second-class vertices, offline vertices with different $y_{j}$ will have different match rates. More importantly, the match rates will be different from the offline vertices in $G$.
There is no reason that $y_{j}=1-\ln 2$ is not the worst case. The idea to solve this issue is that, we can manually force some second-class edges to be redefined as first-class edges. While these edges are still incident to second-class vertices, we can make them act as first-class edges by not allowing them to steal the other half of the arrival rate of the incident online vertex when the other offline neighbor has been matched, as we will see later. So, for each offline vertex $j$ such that $y_{j}<1-\ln 2$, we pick some incident second-class edges with a total matching fraction of $1-\ln 2-y_{j}$ [^6] Of course, we may have to pick “a part of” an edge to guarantee that the total fraction is exactly that value. This is not a problem, since an online vertex type of arrival rate $\lambda$ is equivalent to two online types of arrival rates $\lambda_{1}$ and $\lambda_{2}$ with the same offline neighbors, where $\lambda_{1}+\lambda_{2}=\lambda$. So we only need to split one online type into two for each offline vertex., and redefine them as first-class edges.
Now, a first-class edge is possible to be incident to a second-class vertex, but we successfully guarantee that each offline vertex is matched to first-class edges with exactly $1-\ln 2$ fraction.
The last thing before seeing the generalized algorithm is to define some important notations. Let $g(t)$ denote the probability that, when running Algorithm 3 on $G$, both offline vertices have been matched at time $t$. By symmetry, the two offline vertices are matched at the same rate. For any of them, let $f(t)$ denote the probability that it has been matched at time $t$. Let $\bar{g}(t)=1-g(t)$, which is the probability that at least one offline vertex is unmatched at time $t$. Similarly let $\bar{f}(t)=1-f(t)$.
When applying our generalized algorithm on any instance, we will use $f_{u}(t)$ to denote the probability that the offline vertex $u$ has been matched at time $t$, and $g_{u,v}(t)$ to denote the probability that both $u$ and $v$ have been matched at time $t$. Let $\bar{f}_{u}(t)=1-f_{u}(t)$ and $\bar{g}_{u,v}(t)=1-g_{u,v}(t)$.
Algorithm 4 Generalized version of Algorithm 3
1:When a first-class online vertex arrives, match it to the only neighbor (if possible).
2:When a second-class online vertex arrives, match it along each first-class incident edge (if possible) with probability $\frac{1}{2}$.
3:When a second-class online vertex arrives at time $t>t_{0}$, match it along each second-class incident edge (if possible) with probability $\frac{\bar{g}(t)}{2\bar{g}_{u,v}(t)}$ if the other offline neighbor is unmatched, or with probability $\frac{\bar{g}(t)}{\bar{g}_{u,v}(t)}$ if the other offline neighbor is matched.
Before proving that Algorithm 4 simulates Algorithm 3 well, we need to make sure that for each second-class online vertex, the total probability we assigned to its incident edges does not exceed $1$. This can be guaranteed by the following claim.
Claim 6.
When running Algorithm 4, the probability $\frac{\bar{g}(t)}{\bar{g}_{u,v}(t)}$ is always at most $1$.
To prove 6, we need an auxiliary algorithm. Let $g^{\prime}_{u,v}(t)$ denote the probability that both $u$ and $v$ are unmatched at time $t$. Let $g^{\prime}(t)$ denote the probability that, when running Algorithm 3 on $G$, both offline vertices are unmatched at time $t$. Note that $\bar{g}(t)=2\bar{f}(t)-g^{\prime}(t)$.
Algorithm 5 Auxiliary version of Algorithm 4
1:When a first-class online vertex arrives, match it to the only neighbor (if possible).
2:When a second-class online vertex arrives, match it along each first-class incident edge (if possible) with probability $\frac{1}{2}$.
3:When a second-class online vertex arrives at time $t>t_{0}$, match it along each second-class incident edge (if possible) with probability $\frac{1}{2}\min\{\frac{\bar{g}(t)}{2\bar{f}_{u}(t)-g^{\prime}_{u,v}(t)},1\}$ if the other offline neighbor is unmatched, or with probability $\min\{\frac{\bar{g}(t)}{2\bar{f}_{u}(t)-g^{\prime}_{u,v}(t)},1\}$ if the other offline neighbor is matched.
It’s clear that Algorithm 5 is a valid algorithm. We will show that, when running Algorithm 5, we always have $2\bar{f}_{u}(t)-g^{\prime}_{u,v}(t)\geq\bar{g}(t)$ and $2\bar{f}_{u}(t)-g^{\prime}_{u,v}(t)=\bar{g}_{u,v}(t)$. This means Algorithm 5 is equivalent to Algorithm 4, and proves 6.
Lemma 7.
When running Algorithm 5, we always have $f_{u}(t)\leq f(t)$.
Proof.
Clearly, $f_{u}(t)=f(t)$ when $t\leq t_{0}$. So we start with $f_{u}(t_{0})=f(t_{0})$ and only consider $t>t_{0}$. For each online neighbor $i$ of $u$, let $p_{i,u}(t)$ denote the probability that, when an online vertex of type $i$ arrives at time $t$, it is matched with $u$. By definition,
$\frac{\mathrm{d}}{\mathrm{d}t}f_{u}(t)=\sum_{i}\lambda_{i}p_{i,u}(t).$
By the definition of Algorithm 4, it’s easy to see that $p_{i,u}(t)=\bar{f}_{u}(t)$ for each first-class neighbor $i$ and $p_{i,u}(t)=\frac{1}{2}\bar{f}_{u}(t)$ for each second-class neighbor $i$ such that $(i,u)$ is a first-class edge. So, for each first-class edge $(i,u)$, we have $\lambda_{i}p_{i,u}(t)=x_{iu}\bar{f}_{u}(t)$.
On the other hand, consider any second-class edge $(i,u)$. Let the other neighbor of $i$ be $v$. Since the probability that $u$ is unmatched but $v$ is matched at time $t$ is $\bar{f}_{u}(t)-g^{\prime}_{u,v}(t)$, we have
$p_{i,u}(t)$ $=g^{\prime}_{u,v}(t)\cdot\frac{1}{2}\min\left\{\frac{\bar{g}(t)}{2\bar{f}_{u}(t)-g^{\prime}_{u,v}(t)},1\right\}+(\bar{f}_{u}(t)-g^{\prime}_{u,v}(t))\min\left\{\frac{\bar{g}(t)}{2\bar{f}_{u}(t)-g^{\prime}_{u,v}(t)},1\right\}$
$=\frac{1}{2}(2\bar{f}_{u}(t)-g^{\prime}_{u,v}(t))\min\left\{\frac{\bar{g}(t)}{2\bar{f}_{u}(t)-g^{\prime}_{u,v}(t)},1\right\}$
$\leq\frac{1}{2}\bar{g}(t),$
so $\lambda_{i}p_{i,u}(t)\leq x_{iu}\bar{g}(t)$. Then we have
$\frac{\mathrm{d}}{\mathrm{d}t}f_{u}(t)\leq(1-\ln 2)\bar{f}_{u}(t)+(\ln 2)\bar{g}(t).$
Similarly, by the definition of Algorithm 3, we have
$\frac{\mathrm{d}}{\mathrm{d}t}f(t)$ $=(1-\ln 2)\bar{f}(t)+(2\ln 2)\left(g^{\prime}(t)\cdot\frac{1}{2}+(\bar{f}(t)-g^{\prime}(t))\cdot 1\right)$
$=(1-\ln 2)\bar{f}(t)+(\ln 2)(2\bar{f}(t)-g^{\prime}(t))$
$=(1-\ln 2)\bar{f}(t)+(\ln 2)\bar{g}(t).$
So:
$\frac{\mathrm{d}}{\mathrm{d}t}\left(f(t)-f_{u}(t)\right)$ $\geq(1-\ln 2)(\bar{f}(t)-\bar{f}_{u}(t))$
$=-(1-\ln 2)(f(t)-f_{u}(t)).$
Recall that $f(t_{0})-f_{u}(t_{0})=0$, so $f(t)-f_{u}(t)\geq 0$ for $t\geq t_{0}$. ∎
The above proof also gives us the following corollary.
Corollary 8.
When running Algorithm 5, if we have $2\bar{f}_{u}(t)-g^{\prime}_{u,v}(t)\geq\bar{g}(t)$ for any $t\leq t^{\prime}$, then $f_{u}(t^{\prime})=f(t^{\prime})$.
Proof.
Just notice that, the inequalities in the proof of Lemma 7 become equalities when $2\bar{f}_{u}(t)-g^{\prime}_{u,v}(t)\geq\bar{g}(t)$. ∎
Lemma 9.
When running Algorithm 5, we always have $g_{u,v}(t)\leq g(t)$.
Proof.
Clearly, $g_{u,v}(t)=g(t)$ when $t\leq t_{0}$. So we start with $g_{u,v}(t_{0})=g(t_{0})$ and only consider $t>t_{0}$. When running Algorithm 3, the probability that exactly one offline vertex is unmatched at time $t$ is $2f(t)-g(t)$. In this case, the other one will be matched by both neighbors. So
$\frac{\mathrm{d}}{\mathrm{d}t}g(t)=(2f(t)-g(t))\cdot(1+\ln 2).$
Similarly, when running Algorithm 5, the probability that exactly one of $u$ and $v$ is unmatched at time $t$ is $f_{u}(t)+f_{v}(t)-g_{uv}(t)$. By Lemma 7, this is at most $2f(t)-g_{uv}(t)$. Also notice that the total arrival rate of the neighbors of any offline vertex is $1+\ln 2$. So
$\frac{\mathrm{d}}{\mathrm{d}t}g_{uv}(t)\leq(2f(t)-g_{uv}(t))\cdot(1+\ln 2).$
So we have:
$\frac{\mathrm{d}}{\mathrm{d}t}\left(g(t)-g_{uv}(t)\right)\geq-(1+\ln 2)(g(t)-g_{uv}(t)).$
Recall that $g(t_{0})-g_{uv}(t_{0})=0$, so $g(t)-g_{uv}(t)\geq 0$ for $t\geq t_{0}$. ∎
Lemma 10.
When running Algorithm 5, we always have $2\bar{f}_{u}(t)-g^{\prime}_{u,v}(t)\geq\bar{g}(t)$.
Proof.
Clearly, $2\bar{f}_{u}(t)-g^{\prime}_{u,v}(t)=\bar{g}(t)$ when $t\leq t_{0}$. So we start with $2\bar{f}_{u}(t_{0})-g^{\prime}_{u,v}(t_{0})=\bar{g}(t)$ and only consider $t>t_{0}$. By definition, we have $\bar{g}_{u,v}(t)=\bar{f}_{u}(t)+\bar{f}_{v}(t)-g^{\prime}_{u,v}(t)$. Then, by Lemma 9, we have $\bar{f}_{u}(t)+\bar{f}_{v}(t)-g^{\prime}_{u,v}(t)\geq\bar{g}(t)$. It means that, for any $t$, either $2\bar{f}_{u}(t)-g^{\prime}_{u,v}(t)\geq\bar{g}(t)$, or $2\bar{f}_{v}(t)-g^{\prime}_{u,v}(t)\geq\bar{g}(t)$, or we have both.
Notice that the above functions and their derivatives are all continuous in $[t_{0},1]$. Let $t^{\prime}\in[t_{0},1]$ be the latest time such that for any $t\in[0,t^{\prime}]$ we have both $2\bar{f}_{u}(t)-g^{\prime}_{u,v}(t)\geq\bar{g}(t)$ and $2\bar{f}_{v}(t)-g^{\prime}_{u,v}(t)\geq\bar{g}(t)$. It suffices to show a contradiction for $t^{\prime}<1$. If $t^{\prime}<1$, let $\varepsilon$ be a sufficiently small constant such that, without loss of generality, $2\bar{f}_{u}(t)-g^{\prime}_{u,v}(t)<\bar{g}(t)$ and $2\bar{f}_{v}(t)-g^{\prime}_{u,v}(t)\geq\bar{g}(t)$ for $t\in(t^{\prime},t^{\prime}+\varepsilon]$. Then, since $2\bar{f}_{v}(t)-g^{\prime}_{u,v}(t)\geq\bar{g}(t)$ for $t\leq t^{\prime}+\varepsilon$, we have $f_{v}(t)=f(t)$ for $t\leq t^{\prime}+\varepsilon$ by Corollary 8. Combining with Lemma 7, we get $\bar{f}_{u}(t)\geq\bar{f}_{v}(t)$ for $t\leq t^{\prime}+\varepsilon$. Then we have:
$2\bar{f}_{u}(t)-g^{\prime}_{u,v}(t)\geq\bar{f}_{u}(t)+\bar{f}_{v}(t)-g^{\prime}_{u,v}(t)\geq\bar{g}(t)$
for $t\leq t^{\prime}+\varepsilon$, which is a contradiction. ∎
Combining Lemmas 10 and 8 directly gives us the following corollary.
Corollary 11.
When running Algorithm 5, we always have $f_{u}(t)=f(t)$.
This then gives us the following corollary.
Corollary 12.
When running Algorithm 5, we always have $2\bar{f}_{u}(t)-g^{\prime}_{u,v}(t)=\bar{g}_{u,v}(t)$.
Proof.
By definition, $\bar{g}_{u,v}(t)=\bar{f}_{u}(t)+\bar{f}_{v}(t)-g^{\prime}_{u,v}(t)$. By Corollary 11, $\bar{f}_{u}(t)=\bar{f}_{v}(t)=f(t)$. ∎
As we said before, combining Lemmas 10 and 12 proves 6 and also shows that Algorithm 5 is equivalent to Algorithm 4. Therefore, the above conclusions (in particular, Corollary 8) also work for Algorithm 4. We summarize this in the following lemma.
Lemma 13.
Algorithm 4 is a valid algorithm. Furthermore, when running Algorithm 4, we always have $f_{u}(t)=f(t)$.
We are actually quite close to the end. It only remains to calculate the match rate of each edge from Lemma 13. For each edge $(i,u)$, let $q_{i,u}(t)$ denote the probability that $(i,u)$ has been matched at time $t$. Let $q_{1}(t)$ (resp. $q_{2}(t)$) denote the probability that, when running Algorithm 3 on $G$, a particular first-class (resp. second-class) edge has been matched at time $t$.
Lemma 14.
When running Algorithm 4, for any first-class edge $(i,u)$, we always have $\frac{q_{i,u}(t)}{x_{iu}}=\frac{q_{1}(t)}{1-\ln 2}$; for any second-class edge $(i,u)$, we always have $\frac{q_{i,u}(t)}{x_{iu}}=\frac{q_{2}(t)}{\ln 2}$.
Proof.
Consider any first-class edge $(i,u)$ such that $i$ is first-class, which means $\lambda_{i}=x_{iu}$. At time $t$, the probability that $u$ is unmatched is $\bar{f}_{u}(t)$, which equals $\bar{f}(t)$ by Lemma 13. When $u$ is unmatched and $i$ arrives, we will match $(i,u)$ with probability $1$. So we have
$\frac{\mathrm{d}}{\mathrm{d}t}q_{i,u}(t)=\bar{f}(t)\cdot\lambda_{i}\cdot 1=x_{iu}\bar{f}(t).$
Similarly,
$\frac{\mathrm{d}}{\mathrm{d}t}q_{1}(t)=(1-\ln 2)\bar{f}(t).$
In addition, they all start with $q_{i,u}(0)=q_{1}(0)=0$. Therefore $\frac{q_{i,u}(t)}{x_{iu}}=\frac{q_{1}(t)}{1-\ln 2}$.
Consider any first-class edge $(i,u)$ such that $i$ is second-class, which means $\lambda_{i}=2x_{iu}$. When $u$ is unmatched and $i$ arrives, we will match $(i,u)$ with probability $1/2$. So we have
$\frac{\mathrm{d}}{\mathrm{d}t}q_{i,u}(t)=\bar{f}(t)\cdot\lambda_{i}\cdot\frac{1}{2}=x_{iu}\bar{f}(t).$
So we still have $\frac{q_{i,u}(t)}{x_{iu}}=\frac{q_{1}(t)}{1-\ln 2}$.
Now, consider any second-class edge $(i,u)$, where $i$ must be second-class. Let the other neighbor of $i$ be $v$. At time $t$, the probability that both $u$ and $v$ are unmatched is $g^{\prime}_{u,v}(t)$. In this case, when $i$ arrives, we will match $(i,u)$ with probability $\frac{\bar{g}(t)}{2\bar{g}_{u,v}(t)}=\frac{\bar{g}(t)}{2(2\bar{f}_{u}(t)-g^{\prime}_{u,v}(t))}$. The probability that $u$ is unmatched but $v$ has been matched is $\bar{f}_{u}(t)-g^{\prime}_{u,v}(t)$. In this case, when $i$ arrives, we will match $(i,u)$ with probability $\frac{\bar{g}(t)}{2\bar{f}_{u}(t)-g^{\prime}_{u,v}(t)}$. So we have
$\frac{\mathrm{d}}{\mathrm{d}t}q_{i,u}(t)$ $=g^{\prime}_{u,v}(t)\cdot\lambda_{i}\cdot\frac{\bar{g}(t)}{2(2\bar{f}_{u}(t)-g^{\prime}_{u,v}(t))}+(\bar{f}_{u}(t)-g^{\prime}_{u,v}(t))\cdot\lambda_{i}\cdot\frac{\bar{g}(t)}{2\bar{f}_{u}(t)-g^{\prime}_{u,v}(t)}$
$=x_{i,u}\cdot\frac{\bar{g}(t)}{2\bar{f}_{u}(t)-g^{\prime}_{u,v}(t)}\cdot(g^{\prime}_{u,v}(t)+2(\bar{f}_{u}(t)-g^{\prime}_{u,v}(t)))$
$=x_{i,u}\bar{g}(t).$
Similarly,
$\frac{\mathrm{d}}{\mathrm{d}t}q_{2}(t)$ $=g^{\prime}(t)\cdot(2\ln 2)\cdot\frac{1}{2}+(\bar{f}(t)-g^{\prime}(t))\cdot(2\ln 2)\cdot 1$
$=(\ln 2)(g^{\prime}(t)+2(\bar{f}(t)-g^{\prime}(t)))$
$=(\ln 2)\bar{g}(t).$
In addition, they all start with $q_{i,u}(0)=q_{2}(0)=0$. Therefore $\frac{q_{i,u}(t)}{x_{iu}}=\frac{q_{2}(t)}{\ln 2}$. ∎
Theorem 15.
Algorithm 4 is $\Gamma$-competitive against the optimal value of the Jaillet-Lu LP, where $\Gamma\approx 0.66217>0.662$.
Proof.
Lemma 5 tells us that $\frac{q_{1}(t)}{1-\ln 2}=\frac{q_{2}(t)}{\ln 2}=\Gamma\approx 0.66217$. Combining it with Lemma 14, we know that every edge $(i,j)$ is matched with probability $\Gamma x_{ij}$. Then the expected objective of Algorithm 4 is $\Gamma\sum_{(i,j)\in E}w_{ij}x_{ij}$. Comparing it with the optimal LP value $\sum_{(i,j)\in E}w_{ij}x_{ij}$ proves the theorem. ∎
Finally, we remark that, to implement Algorithm 4, we need to compute $\bar{g}(t)$ and $\bar{g}_{u,v}(t)$ efficiently. In Section A.1, we have the closed-form expression of $g(t)$ which can be computed in $O(1)$ time. However, $\bar{g}_{u,v}(t)$ is related to the instance itself, and it might take exponential time for the algorithm to compute it precisely. Nevertheless, we can use the standard Monte-Carlo method to estimate it with an $o(1)$ multiplicative error in polynomial time. It then gives us a polynomial-time $(\Gamma-o(1))$-competitive algorithm. See Section A.3 for more details.
Acknowledgments
We thank Mikkel Thorup for helpful comments.
Appendix A Appendix
A.1 Missing Details in Theorem 4
The expected objective of Algorithm 2 can be computed as follows. Let $f(t)$ denote the probability that exactly one offline vertex has been matched at time $t$ and $g(t)$ denote the probability that both offline vertices have been matched at time $t$. By our matching strategy in Algorithm 2, during time $[0,t_{0}]$, each offline vertex will only be matched if its first-class neighbor arrives. So for $t\in[0,t_{0}]$, we have
$f(t)$ $=2\left(1-e^{-(1-\ln 2)t}\right)e^{-(1-\ln 2)t},$
$g(t)$ $=\left(1-e^{-(1-\ln 2)t}\right)^{2}.$
Now consider $t\in[t_{0},t_{1}]$. When one offline vertex has been matched, the other one will only be matched by its first-class neighbor, so
$\frac{\mathrm{d}}{\mathrm{d}t}g(t)=f(t)\cdot(1-\ln 2).$
When both offline vertices are unmatched, as long as any online vertex arrives, one of the offline vertices will be matched. Note that the probability that both offline vertices are unmatched at time $t$ is simply $(1-f(t)-g(t))$. So
$\frac{\mathrm{d}}{\mathrm{d}t}f(t)=(1-f(t)-g(t))\cdot 2-\frac{\mathrm{d}}{\mathrm{d}t}g(t).$
Combining these differential equations with the initial values $f(t_{0})$ and $g(t_{0})$, we get, for $t\in[t_{0},t_{1}]$,
$f(t)=2^{t+1}e^{-t}-\frac{2^{2t_{0}+1}e^{-2t}+2^{t+t_{0}+1}e^{-t-t_{0}}\ln 2}{1+\ln 2}$
and
$g(t)=1-2^{t+1}e^{-t}+\frac{2^{2t_{0}}e^{-2t}(1-\ln 2)+2^{t+t_{0}+1}e^{-t-t_{0}}\ln 2}{1+\ln 2}.$
Now consider $t\in[t_{1},1]$. When one offline vertex has been matched, the other one will be matched by both neighbors, so
$\frac{\mathrm{d}}{\mathrm{d}t}g(t)=f(t)\cdot(1+\ln 2).$
And we still have
$\frac{\mathrm{d}}{\mathrm{d}t}f(t)=(1-f(t)-g(t))\cdot 2-\frac{\mathrm{d}}{\mathrm{d}t}g(t).$
Combining these differential equations with the initial values $f(t_{1})$ and $g(t_{1})$, we get, for $t\in[t_{1},1]$,
$f(t)=2^{2t_{1}+1-t}e^{-t}-\frac{2^{2t_{0}+1}e^{-2t}}{1-\ln 2}-\frac{2^{2t_{1}+t_{0}+1}e^{-t-t_{0}}\ln 2}{2^{t}(1+\ln 2)}+\frac{2^{2t_{0}+t_{1}+2}e^{-t-t_{1}}\ln 2}{2^{t}(1+\ln 2)(1-\ln 2)}$
and
$g(t)=1-2^{2t_{1}+1-t}e^{-t}+\frac{2^{2t_{0}}e^{-2t}(1+\ln 2)}{1-\ln 2}+\frac{2^{2t_{1}+t_{0}+1}e^{-t-t_{0}}\ln 2}{2^{t}(1+\ln 2)}-\frac{2^{2t_{0}+t_{1}+2}e^{-t-t_{1}}\ln 2}{2^{t}(1+\ln 2)(1-\ln 2)}.$
Then we know that the expected number of offline vertices we have matched is $f(1)+2g(1)$. To compute the expected objective, it remains to compute how many first-class edges we have matched. Let $p(t)$ denote the expected number of first-class edges we have matched at time $t$. We have
$\frac{\mathrm{d}}{\mathrm{d}t}p(t)=(1-f(t)-g(t))\cdot 2(1-\ln 2)+f(t)\cdot(1-\ln 2).$
Solving it, we get
$p(1)=2+2^{2t_{0}}e^{-2}\ln 2-\frac{2^{2t_{1}}e^{-1}(1-\ln 2)}{1+\ln 2}+\frac{P}{(1+\ln 2)^{2}},$
where
$P=$ $2\displaystyle\ 2^{t_{0}+2t_{1}}e^{-t_{0}-1}(1-\ln 2)\ln 2+2^{2t_{0}+1}e^{-2t_{1}}(1-\ln 2)\ln 2+2^{t_{0}+t_{1}+2}e^{-t_{0}-t_{1}}\ln^{2}$ 2
$-2^{2t_{0}+t_{1}+1}e^{-t_{1}-1}\ln 2-2^{t_{1}+2}e^{-t_{1}}(1+\ln 2)\ln 2-2^{2t_{0}}e^{-2t_{0}}(1-\ln 2)^{2}\ln 2.$
Then, the expected objective is $\textsc{Alg}(k,t_{0},t_{1})=f(1)+2g(1)+(k-1)p(1)$.
A.2 Missing Details in Lemma 5
Given Section A.1, it’s straightforward to compute all we need. In the end, each first-class edge is matched with probability $q_{1}=p(1)/2$. Each second-class edge is matched with probability $q_{2}=(f(1)+2g(2)-p(1))/2$. Forcing $t_{1}=t_{0}$ and balancing $q_{1}/(1-\ln 2)=q_{2}/\ln 2$, we get $q_{1}/(1-\ln 2)=q_{2}/\ln 2\approx 0.66217$ when $t_{0}\approx 0.14753$.
A.3 Computational Aspects of Algorithm 4
The matching strategy of Algorithm 4 depends on which offline vertices are still unmatched. If we want to compute $\bar{g}_{u,v}(t)$ precisely, one way is to define a function $f_{S}(t)$ for each set $S\subseteq J$, and solve the differential equations between them as we did in Section A.1. However, there will be exponentially many equations, so we do not expect to solve them efficiently.
Here we discuss how to efficiently estimate $\bar{g}_{u,v}(t)$ with an $o(1)$ multiplicative error. First, notice that $\bar{g}_{u,v}(t)=\Omega(1)$ for $t\geq t_{0}$, which means we only need to achieve an $o(1)$ additive error. Since $\bar{g}_{u,v}(t)$ is a continuous function with bounded derivative, we can approximate $t$ by its closest multiple of $1/poly(\Lambda)$. Therefore, we only need to consider polynomially many discrete time points, which means the total number of $\bar{g}_{u,v}(t)$’s we need to estimate is also polynomial. Then we can use the standard Monte-Carlo method to estimate them in polynomial time, with $o(1)$ additive errors.